# Hypergraph Characterization of Fusion Rings

### v1

@@ 5. Low Rank Fusion Rings
$2$ $(1,2)$, $(1,5)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[21, 20]
$4$ $(1,3)$, $(1,4)$, $(1,5)$, $(2,5)$, $(3,4)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{Ising}$[21]
$2$ $(1,2)$, $(1,4)$, $(2,4)$, $(3,2)$, $(4,2)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}\left(3\right)_{2}$[21]
$1$ $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
$2$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,3)$, $(2,4)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$
$1$ $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
$1$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$, $(5,2)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$
$1$, $2$ $(1,2)$, $(1,5)$, $(2,1)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$
$1$, $2$ $(1,2)$, $(1,4)$, $(2,1)$, $(2,4)$, $(3,1)$, $(3,2)$, $(4,1)$, $(4,2)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
$1$, $2$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$
$1$, $2$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,5)$, $(3,4)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[20]
$1$, $3$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,5)$, $(5,3)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(2,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(3)_{2}$ [20]
$1$, $2$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,2)$, $(3,4)$, $(4,2)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(3,4,5)$
$1$, $2$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,1)$, $(3,2)$, $(3,4)$, $(4,2)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(2,3,5)$, $(3,4,5)$
$1$, $2$, $3$, $4$ $(1,5)$, $(2,5)$, $(3,5)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$
$1$, $2$, $3$, $4$ $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$
$1$, $2$, $3$, $4$ $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,4)$, $(4,1)$, $(4,2)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
$1$, $2$, $4$, $5$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(2,4,5)$

### v2

@@ 11 $(1,2)$, $(1,3)$, $(1,4)$ $(2,3,4)$ not categorifiable[11]
22 $(1,2)$, $(1,5)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[24, 23]
44 $(1,3)$, $(1,4)$, $(1,5)$, $(2,5)$, $(3,4)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{Ising}$[24]
22 $(1,2)$, $(1,4)$, $(2,4)$, $(3,2)$, $(4,2)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}\left(3\right)_{2}$[24]
11 $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
22 $(1,2)$, $(1,3)$, $(1,4)$, $(2,3)$, $(2,4)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$
11 $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
11 $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$, $(5,2)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$
11, 22 $(1,2)$, $(1,5)$, $(2,1)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$
11, 22 $(1,2)$, $(1,4)$, $(2,1)$, $(2,4)$, $(3,1)$, $(3,2)$, $(4,1)$, $(4,2)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
11, 22 $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$
11, 22, 44 $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,5)$, $(3,4)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[23]
11, 33, 44 $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,5)$, $(5,3)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(2,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(3)_{2}$ [23]
11, 22, 44 $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,2)$, $(3,4)$, $(4,2)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(3,4,5)$
11, 22, 44 $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,1)$, $(3,2)$, $(3,4)$, $(4,2)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(2,3,5)$, $(3,4,5)$
11, 22, 33, 44 $(1,5)$, $(2,5)$, $(3,5)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$
11, 22, 33, 44 $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$
11, 22, 33, 44 $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,4)$, $(4,1)$, $(4,2)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
11, 22, 44, 55 $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(2,4,5)$